40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 20 00:59:02 2011, kiwamu, Summary, General, Weekly report 
    Reply  Mon Jun 27 10:04:29 2011, kiwamu, Summary, General, Weekly report 
       Reply  Mon Jul 4 14:27:35 2011, kiwamu, Summary, General, Weekly report 
          Reply  Mon Jul 25 11:02:19 2011, kiwamu, Summary, General, Weekly report 
             Reply  Mon Jul 25 11:04:22 2011, kiwamu, Summary, General, Weekly report 
                Reply  Mon Jul 25 11:06:38 2011, kiwamu, Summary, General, Weekly report 
                   Reply  Mon Aug 1 22:30:59 2011, kiwamu, Summary, General, Weekly report 
Message ID: 4936     Entry time: Mon Jul 4 14:27:35 2011     In reply to: 4890     Reply to this: 5026
Author: kiwamu 
Type: Summary 
Category: General 
Subject: Weekly report 

Summary of the week ending July 3rd.  Number of elog entries = 44 

- SUS

   * The output TO_COIL matrix were simplified
   * Checked all the BO whitening switch => Only ITMY_UL didn't switch
   * All the DOF filters were normalized.
       => All the DOF filters are now ("3:0.0", "Cheby", "BounceRoll") 
       => The High pass should have 30Hz cut off ("30:0.0") ?
   * All the resonant peaks has been fit

- LSC

   * MICH noise budget.
               => dominated by sensing noise.
   * The sensing matrix in the PRMI configuration was measured. 
               => The demodulation phase on AS55 seemed wrong. Need a doublecheck
   * A new screen, called C1LSC_OVERVIEW.adl, was released.

   * A channel name modification: "PRC" and "SRC" => "PRCL" and "SRCL" and etc.
   * The response of the LSC whitening filters were checked. 
                 => CH26 showed different phase response.
 

- MC work

 * Power budget on the AP table was made (in a high power situation).
REFL11 = 7.4 mW
REFL55 = 22 mW
        MCREFL = 114 mW
        WFS1   = 1.24 mW
        WFS2   = 2.7 mW
 * Measurement and adjustment of RFQPD response
         Resonance frequencies of WFS1 and WFS2 were adjusted. WFS1 and WFS2 were installed on the AP table
 * Started working on MCL path. 
         => needs some more CDS jobs to correctly assign ADC channels

- CDS

 * Joe modified the automated scripts for producing model webviews

- ABSL

  * The alignment of the injection beam was done.

- Fiber experiment

 * A fiber was laid down from the ETMY table to the PSL table

- TT characterization

 * The mechanical stage for the horizontal displacement measurements is set up. 

- Configuration and other topics

      * Maglev stuff has gone to bridge lab.
      * Chris.W told us that the EPICS mutex issue can be solved by upgrading the EPICS version
      * All the PDs are stored in the east arm cabinet E4
      * Safety interlocks were connected to the ETMY laser and ABSL laser
      * Cshrc.40m was modified to make 32-bit machine happy
      * NDS2 buffer size on Mafalada had been too small and was increased somewhat such that we can still work for the SUS peak fit job
 
ELOG V3.1.3-